OpenStudy (bookworm14):

last 2 (sry went back over to any i had trouble with) @satellite73

OpenStudy (bookworm14):

I got B, but it was a guess

OpenStudy (anonymous):

\[\frac{x}{4}=\frac{11}{5}\] \[x=\frac{4\times 11}{5}\] whatever that is

OpenStudy (anonymous):

yeah, B

OpenStudy (bookworm14):

this one I could not figure out so I guessed also I chose B

OpenStudy (anonymous):

\[\frac{13}{8}=\frac{x}{7}\] making \[x=\frac{7\times 13}{8}\] then a calculator

OpenStudy (bookworm14):

so I got it right !?

OpenStudy (anonymous):

did you get 11.375?

OpenStudy (anonymous):

oh 11.4 ok

OpenStudy (anonymous):

that it?

OpenStudy (bookworm14):

i think so yes

OpenStudy (bookworm14):

im gonna go turn it in fwew, thank you for all your help !!!

OpenStudy (anonymous):

yw bet you get 100!